### v1

@@ 6.1 Trade dynamic networks
We use the trade dynamic networks between countries for different food and agriculture products during the year from $2005$ to $2022$. The data is collected by the Food and Agriculture Organization of the United Nations and is available at https://www.fao.org/faostat/en/#data/TM. We construct a collection of dynamic networks for $T=2022-2005+1=18$ time points, one dynamic network for each food or agriculture product, where vertices represent countries and edges in each network represent trade relationships between countries. For each product $i$ and each time point $t$, we obtain the adjacency matrix $\mathbf{A}^{(i)}_{t}$ by (1) we set $(\mathbf{A}^{(i)}_{t})_{r,s}=(\mathbf{A}^{(i)}_{t})_{s,r}=1$ if there is product $i$ trade between countries $r$ and $s$; (2) we ignore the links between countries $r$ and $s$ in $\mathbf{A}^{(i)}_{t}$ if their total trade amount for the product $i$ at the time point $t$ is less than two hundred thousands US dollars; (3) finally we extract the intersection of the largest connected components of $\{\mathbf{A}^{(i)}_{t}\}$ to get the networks for the common involved countries. The resulting adjacency matrices $\{\mathbf{A}^{(i)}_{t}\}_{i\in[m],t\in[T]}$ corresponding to $m=18$ dynamic networks for $T=18$ time points on a set of $n=58$ vertices.
(a) Hierarchical clustering result
(b) Embeddings based on $\mathbf{D}^{\star}$
Figure 4: Clustering results for product trading patterns
@@
To analyze the relationships between trade patterns of these $m=18$ products, we apply Algorithm 1 to obtain the hierarchical clustering result for the $m=18$ dynamic networks with the embedding dimensions $d$ and $r$ chosen to be $2$. The left panel of Fig. 4 presents the dendrogram of the hierarchical clustering result and we use CMDS to visualize the distance relationships between products in $\mathbf{D}^{\star}$; see the right panel of Fig. 4. From Fig. 4 we see there is a high degree of correlation between the trade pattern relationships and the types of products. More specifically, beer and spirits both are alcoholic beverages produced from grain-based materials with similar consumer bases, and they show very similar trade patterns in the figure. Crude organic material n.e.c. and food preparations n.e.c. both have a certain ambiguity in their classification, and their trade trends are also be more challenging to match with other clearly classified products. Except the above four products, there appears to be two main clusters formed by raw/unprocessed products (top cluster) and processed products (bottom cluster), and it suggests discernable differences in the trading patterns for these types of products.
6.2 Drosophila larval connectome data
The complete synaptic-resolution connectome of the Drosophila larval brain was recently completed [55], enabling the generation of biologically realistic models of its neural circuit based on known anatomical connectivity [19]. Training such realistic models in simulations provides significant computational power and flexibility, enabling us to study and enhance our understanding of how real animal brains operate. This approach has been applied in recent papers [1] to investigate the operational mechanisms of the Drosophila larval brain.
The connectivity-constrained model proposed in [19, 28] can be used to generate a recurrent network of the larval mushroom body with feedback neurons, where the connectome can be constrained, and in this process, a series of stimuli can be delivered to the network; see [19, 28] for more details. Therefore, by analyzing the dynamic networks generated by this model after the removal of particular edges between neurons, we can indirectly study the function of these edges.
In this experiment, we consider the process of learning a new association between a conditioned stimulus (CS) (for example, an odor) and rewards or punishments (unconditioned stimuli, US), which serve as reinforcement. We simulate the activity of the dynamic neuron network for a total of $T=160$ time points, and at $t=16$, a random odor is delivered to the neurons in the mushroom body (CS1), followed by a reward delivered at $t=20$. We set all stimuli to last for $3$ time points. After this initial CS of odor with the reward, the odor is also delivered again without the reward at $t=80$ (CS2) and $t=140$ (CS3). The association may be weakened by exposure to the same CS without reinforcement during CS2, so we may observe a decrease in network attraction from CS2 to CS3. Here, we define network attraction as the ratio of the strength of activity in neurons responsible for attraction to the strength of activity in neurons responsible for aversion. Thus, this process includes the learning of the association and its extinction. A schematic of this process can be seen in Fig. 5.
Figure 5: Schematic of the process including the learning of a association and its extinction.
We applied the DNCMD algorithm to obtain the dendrogram shown in Fig. 6, with the embedding dimension for vertex latent positions set to $d=5$ and the dimension for mirror set to $r=3$. In Fig. 6, we can clearly see that most of the dynamic networks resulting from the removal of edge $6$ form a distinct cluster, indicating that edge $6$ plays a critical and unique role in this process.
Figure 6: Dendrogram of dynamic networks of neurons for the removal of each single edge, obtained using the DNCMD algorithm. The labels represent the index of the removed edge (from $1$ to $13$) and are shown in different colors, as indicated in the legend on the right, and for each removed edge we have $11$ replicates.
@@ 6.2 Drosophila larval connectome data
Based on the dendrogram in Fig. 6, we develop a numerical measure to quantify the degree of concentration of replicates for each label. Given a number of clusters $K$, we can obtain a clustering result from the dendrogram and use a contingency table to display the frequency distribution of the clusters across the labels; see Panel (a) of Fig. 7 for the contingency table when $K=3$ is set. We denote the frequencies in the contingency table for $K$ clusters by $\{\mathbf{C}_{i,k}^{(K)}\}_{i\in[13],k\in[K]}$, where $i$ is the index over the $13$ labels, $k$ is the index over the total $K$ clusters. Since each label has $11$ observations, we compute the frequency rate among all dynamic networks with label $i$ as $\mathbf{R}_{i,k}^{(K)}:=\mathbf{C}_{i,k}^{(K)}/11$; see Panel (b) of Fig. 7 for an example table of the corresponding frequency rates. For any fixed $K$, if a label $i$ has a high maximum frequency rate $\max_{k\in[K]}\{\mathbf{R}_{i,k}^{(K)}\}$, it suggests that the dynamic networks for this label are concentrated in one of the clusters. Instead of using the maximum frequency rates for a specific $K$, we consider integrating them over a reasonable range of $K$. For example, we tune $K$ from $1$ to $K_{\max}=10$ to obtain the curve of the maximum frequency rate for each label, as shown in Panel (a) of Fig. 8. Then, we compute the area under the curve (AUC) for each label, and normalize it by dividing the AUC by $(K_{\max}-1)$, scaling it between 00 and $1$; see Panel (b) of Fig. 8 for the normalized AUCs for all labels. Based on the normalized AUCs, we also see that the replicates for edge $6$ are the most concentrated. Figures 7 and 8 quantitatively support our claim that edge $6$ is special.
@@
Figure 7: Panel (a) shows the contingency table for the case of $K=3$ clusters. Panel (b) shows the corresponding frequency rates, and the maximum frequency rate for each label is highlighted in yellow.
@@
Figure 8: Panel (a) shows the maximum frequency rates for all labels when we tune $K$ from $1$ to $K_{\max}=10$. In panel (a), the two dashed curves respectively describe the maximum value, $1$, and the minimum value, $1/K$, that the maximum frequency rate can achieve. Panel (b) shows the normalized AUCs for all labels, and the two dashed lines respectively describe the maximum value and the minimum value that the normalized AUC can achieve.
@@
One more question we are also concerned with is which edges might have similar functions. To address this, we measure the similarity between the distributions of replicates across clusters for each pair of edges. Given a number of clusters $K$, for each pair of labels, we construct a measure of similarity using the Jaccard distance between their distributions across the clusters. The Jaccard distance ranges from 00 to $1$, where 00 indicates that the two distributions are exactly the same. Thus, a small Jaccard distance suggests that the dynamic networks for these two labels have similar distributions across the clusters. Fig. 9 shows the Jaccard distances for all pairs of labels when $K=3$. We also consider integrating the Jaccard distances over a range of $K$. We show the Jaccard distances for all pairs of labels when we tune $K$ from $1$ to $K_{\max}=10$ in Panel (a) of Fig. 10. We then obtain their AUCs and also normalize it by dividing the AUC by $/(K_{\max}-1)$ in Panel (b) of Fig. 10. A pair of labels $(i,j)$ with a small normalized AUC of the Jaccard distance suggests that this pair of labels always has similar distributions across the clusters, further implying that the dynamic networks corresponding to these two edges have similar distributions. From Panel (b) of Fig. 10, we observe that edges $1$ and $7$, as well as edges $5$ and $9$, exhibit the most similar functions, whereas edge $6$ is distinct from all other edges.
@@
Figure 9: Jaccard distances between the distributions across the clusters for all pair of labels for $K=3$.
(a) Jaccard distance for all pairs of labels when we have $K$ clusters
(b) Normalized AUC for pairs of labels
@@
Figure 10: Panel (a) shows the Jaccard distances for all pairs of labels when we tune $K$ from $1$ to $K_{\max}=10$. The dashed line represents the maximum value, $1$, that the Jaccard distance can achieve. We have total $\binom{13}{2}=78$ lines in Panel (a) to represent all pairs of these $13$ labels in different colors. Panel (b) shows the normalized AUCs for all pairs of labels.
7 Conclusion and Discussion
In summary, this paper presents DNCMD for clustering dynamic networks according to their evolution patterns. Our algorithm is based on the mirror method, which captures the important features of the evolution of dynamic networks by curves in low-dimensional Euclidean space, and uses orthogonal Procrustes analysis to eliminate the non-identifiability of the mirrors, with the resulting pairwise distances used for hierarchical clustering to produce a dendrogram and the final clustering outcome. Under mild assumptions, we establish theoretical guarantees for the exact recovery of our algorithm in two general scenarios: when vertex latent positions are deterministic and when vertex latent positions are randomly generated from stochastic processes. The simulation experiments validate our theoretical results and demonstrate that our algorithm generally provides highly accurate clustering results. We finally demonstrate the applications of our algorithm on real-world data. We cluster trade dynamic networks to analyze the relationships in the evolution of food products. For the Drosophila larval connectome data, we show how to classify the dynamic neural networks obtained after removing each specific edge and leverage this result to indirectly analyze the function of edges in specific neural processes, such as learning a new association between an odor stimulus and a reward.
We now mention potential directions for future work. Our clustering approach is based on mirrors, and the mirror idea can be extended to other network models beyond RDPG. For instance, when the common subspace independent edge (COSIE) model [6] is applied to a dynamic network, the probability matrix at time point $t$ can be written as $\mathbf{P}_{t}=\mathbf{U}\mathbf{R}_{t}\mathbf{U}^{\top}$, where $\mathbf{U}$ represents the invariant latent structure, while the possibly time-varying connectivity pattern is modeled by $\{\mathbf{R}_{t}\}_{t}$. Following the mirror idea, we can measure the pairwise differences between $\{\mathbf{R}_{t}\}_{t}$, which represent the evolving underlying structure, to find the corresponding curve of the dynamic network in low-dimensional space under this model. A feasible measure for the difference between $\hat{\mathbf{R}}_{t_{1}}$ and $\hat{\mathbf{R}}_{t_{2}}$ for any $t_{1}$ and $t_{2}$ can be found in Theorem 6 of [58]. Finally, the mirror approach, which finds low-dimensional configurations to extract main signals, along with the corresponding clustering method based on distances between mirrors, can be applied to problems beyond dynamic networks, as long as the problem involves similar grouped structures.
Appendix A Proof of Theorem 2.5
@@ 7 Conclusion and Discussion
$\mathbf{W}:=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\mathbf{X}_{t_{1}}\mathbf{O}-\mathbf{X}_{t_{2}}\|_{F},\quad\hat{\mathbf{W}}:=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}_{t_{1}}\mathbf{O}-\hat{\mathbf{X}}_{t_{2}}\|_{F},$
$\mathbf{W}^{(t_{1})}:=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}_{t_{1}}\mathbf{O}-\mathbf{X}_{t_{1}}\|_{F},\quad\mathbf{W}^{(t_{2})}:=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}_{t_{2}}\mathbf{O}-\mathbf{X}_{t_{2}}\|_{F}.$
@@
(8) $n^{1/2}|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|$ $=\big{|}\|\hat{\mathbf{X}}_{t_{1}}\hat{\mathbf{W}}-\hat{\mathbf{X}}_{t_{2}}\|_{F}-\|\mathbf{X}_{t_{1}}\mathbf{W}-\mathbf{X}_{t_{2}}\|_{F}\big{|}$
@@
For $\hat{\mathbf{X}}_{t_{1}}\mathbf{W}^{(t_{1})}-\mathbf{X}_{t_{1}}$ and $\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{2}}$, according to Theorem 2.1 in [48] under the assumption $n\rho_{n}=\Omega(\log n)$ we have
(9) $\|\hat{\mathbf{X}}_{t_{1}}\mathbf{W}^{(t_{1})}-\mathbf{X}_{t_{1}}\|_{F}\lesssim 1\text{ and }\|\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{2}}\|_{F}\lesssim 1$
@@
(10) $\|\mathbf{X}_{t_{1}}\|\asymp(n\rho_{n})^{1/2}\text{ and }\|\mathbf{X}_{t_{2}}\|\asymp(n\rho_{n})^{1/2}.$
@@
(11) $\|\mathbf{W}-\mathbf{W}^{(t_{1})\top}\hat{\mathbf{W}}\mathbf{W}^{(t_{2})}\|\leq\frac{2\|\mathbf{W}^{(t_{1})\top}\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}\|}{\sigma_{\min}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})}.$
Indeed, $\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}$ is invertible. Now suppose $\|\mathbf{W}^{(t_{1})\top}\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}\|<\sigma_{\min}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})$. Then $\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}$ is also invertible and hence, by Theorem 1 in [36] we have
@@
Otherwise if $\|\mathbf{W}^{(t_{1})\top}\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}\|\geq\sigma_{\min}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})$ then, as $\|\mathbf{W}-\mathbf{W}^{(t_{1})\top}\hat{\mathbf{W}}\mathbf{W}^{(t_{2})}\|\leq 2$, Eq. (11) holds trivially. By Eq. (9) and Eq. (10) we have
@@
with high probability. Notice we have $\lambda_{k}(\mathbf{P}_{t_{1}})\asymp n\rho_{n}$ and $\lambda_{k}(\mathbf{P}_{t_{2}})\asymp n\rho_{n}$ for any $k\in[d]$, it follows that $\sigma_{k}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})\asymp n\rho_{n}$ for any $k\in[d]$. Then by Eq. (11) we have
(12) $\|\mathbf{W}-\mathbf{W}^{(t_{1})\top}\hat{\mathbf{W}}\mathbf{W}^{(t_{2})}\|\lesssim\frac{(n\rho_{n})^{1/2}}{n\rho_{n}}\lesssim(n\rho_{n})^{-1/2}$
with high probability. Combining Eq. (8), Eq. (9), Eq. (10) and Eq. (12) we have

### v2

@@ 6.2 Trade dynamic networks
We use the trade dynamic networks between countries for different food and agriculture products during the year from $2005$ to $2022$. The data is collected by the Food and Agriculture Organization of the United Nations and is available at https://www.fao.org/faostat/en/#data/TM. We construct a collection of dynamic networks for T=20222005+1=18T=2022-2005+1=18 time points, one dynamic network for each food or agriculture product, where vertices represent countries and edges in each network represent trade relationships between countries. For each product $i$ and each time point $t$, we obtain the adjacency matrix $\mathbf{A}^{(i)}_{t}$ by (1) we set $(\mathbf{A}^{(i)}_{t})_{r,s}=(\mathbf{A}^{(i)}_{t})_{s,r}=1$ if there is product $i$ trade between countries $r$ and $s$; (2) we ignore the links between countries $r$ and $s$ in $\mathbf{A}^{(i)}_{t}$ if their total trade amount for the product $i$ at the time point $t$ is less than two hundred thousands US dollars; (3) finally we extract the intersection of the largest connected components of $\{\mathbf{A}^{(i)}_{t}\}$ to get the networks for the common involved countries. The resulting adjacency matrices $\{\mathbf{A}^{(i)}_{t}\}_{i\in[m],t\in[T]}$ corresponding to $m=18$ dynamic networks for $T=18$ time points on a set of $n=58$ vertices.
@@
To analyze the relationships between trade patterns of these $m=18$ products, we apply Algorithm 1 to obtain the hierarchical clustering result for the $m=18$ dynamic networks with the embedding dimensions $d$ and $r$ chosen to be $2$. Figure 7 presents the dendrogram of the hierarchical clustering result. From Figure 7, we see there is a high degree of correlation between the trade pattern relationships and the types of products. More specifically, beer and spirits both are alcoholic beverages produced from grain-based materials with similar consumer bases, and they show very similar trade patterns in the figure. Crude organic material n.e.c. and food preparations n.e.c. both have a certain ambiguity in their classification, and their trade trends are also be more challenging to match with other clearly classified products. Except the above four products, there appears to be two main clusters formed by raw/unprocessed products (top cluster) and processed products (bottom cluster), and it suggests discernable differences in the trading patterns for these types of products.
Figure 7: Hierarchical clustering result for product trading patterns.
7 Conclusion and Discussion
In summary, this paper presents DNCMD for clustering dynamic networks according to their evolution patterns. Our algorithm is based on the mirror method, which captures the important features of the evolution of dynamic networks by curves in low-dimensional Euclidean space, and uses orthogonal Procrustes analysis to eliminate the non-identifiability of the mirrors, with the resulting pairwise distances used for hierarchical clustering to produce a dendrogram and the final clustering outcome. Under mild assumptions, we establish theoretical guarantees for the exact recovery of our algorithm in two general scenarios: when vertex latent positions are deterministic and when vertex latent positions are randomly generated from stochastic processes. The simulation experiments validate our theoretical results and demonstrate that our algorithm generally provides highly accurate clustering results. We finally demonstrate the applications of our algorithm on real-world data. For the Drosophila larval connectome data, we show how to classify the dynamic neural networks obtained after removing each specific edge and leverage this result to indirectly analyze the function of edges in specific neural processes, such as learning a new association between an odor stimulus and a reward. And we cluster trade dynamic networks to analyze the relationships in the evolution of food products.
We now mention potential directions for future work. Our clustering approach is based on mirrors, and the mirror idea can be extended to other network models beyond RDPG. For instance, when the common subspace independent edge (COSIE) model Arroyo et al. (2021) is applied to a dynamic network, the probability matrix at time point $t$ can be written as $\mathbf{P}_{t}=\mathbf{U}\mathbf{R}_{t}\mathbf{U}^{\top}$, where $\mathbf{U}$ represents the invariant latent structure, while the possibly time-varying connectivity pattern is modeled by $\{\mathbf{R}_{t}\}_{t}$. Following the mirror idea, we can measure the pairwise differences between $\{\mathbf{R}_{t}\}_{t}$, which represent the evolving underlying structure, to find the corresponding curve of the dynamic network in low-dimensional space under this model. A feasible measure for the difference between $\hat{\mathbf{R}}_{t_{1}}$ and $\hat{\mathbf{R}}_{t_{2}}$ for any $t_{1}$ and $t_{2}$ can be found in Theorem 6 of Zheng and Tang (2022). Finally, the mirror approach, which finds low-dimensional configurations to extract main signals, along with the corresponding clustering method based on distances between mirrors, can be applied to problems beyond dynamic networks, as long as the problem involves similar grouped structures.
Supplementary Material for “Dynamic networks clustering via mirror distance”
Appendix A Additional Numerical Results
A.1 Additional analysis for Drosophila larval connectome
@@
Based on the dendrogram in Figure 6, we develop a numerical measure to quantify the degree of concentration of replicates for each label. Given a number of clusters $K$, we can obtain a clustering result from the dendrogram and use a contingency table to display the frequency distribution of the clusters across the labels; see Panel (a) of Figure A.1 for the contingency table when $K=3$ is set. We denote the frequencies in the contingency table for $K$ clusters by $\{\mathbf{C}_{i,k}^{(K)}\}_{i\in[13],k\in[K]}$, where $i$ is the index over the $13$ labels, $k$ is the index over the total $K$ clusters. Since each label has $11$ observations, we compute the frequency rate among all dynamic networks with label $i$ as $\mathbf{R}_{i,k}^{(K)}:=\mathbf{C}_{i,k}^{(K)}/11$; see Panel (b) of Figure A.1 for an example table of the corresponding frequency rates. For any fixed $K$, if a label $i$ has a high maximum frequency rate $\max_{k\in[K]}\{\mathbf{R}_{i,k}^{(K)}\}$, it suggests that the dynamic networks for this label are concentrated in one of the clusters. Instead of using the maximum frequency rates for a specific $K$, we consider integrating them over a reasonable range of $K$. For example, we tune $K$ from $1$ to $K_{\max}=10$ to obtain the curve of the maximum frequency rate for each label, as shown in Panel (a) of Figure A.2. Then, we compute the area under the curve (AUC) for each label, and normalize it by dividing the AUC by $(K_{\max}-1)$, scaling it between 0 and $1$; see Panel (b) of Figure A.2 for the normalized AUCs for all labels. Based on the normalized AUCs, we also see that the replicates for edge $6$ are the most concentrated. Figures A.1 and A.2 quantitatively support our claim that edge $6$ is special.
@@
Figure A.1: Panel (a) shows the contingency table for the case of $K=3$ clusters. Panel (b) shows the corresponding frequency rates, and the maximum frequency rate for each label is highlighted in yellow.
@@
Figure A.2: Panel (a) shows the maximum frequency rates for all labels when we tune $K$ from $1$ to $K_{\max}=10$. In panel (a), the two dashed curves respectively describe the maximum value, 11, and the minimum value, $1/K$, that the maximum frequency rate can achieve. Panel (b) shows the normalized AUCs for all labels, and the two dashed lines respectively describe the maximum value and the minimum value that the normalized AUC can achieve.
@@
One more question we are also concerned with is which edges might have similar functions. To address this, we measure the similarity between the distributions of replicates across clusters for each pair of edges. Given a number of clusters $K$, for each pair of labels, we construct a measure of similarity using the Jaccard distance between their distributions across the clusters. The Jaccard distance ranges from 0 to $1$, where 0 indicates that the two distributions are exactly the same. Thus, a small Jaccard distance suggests that the dynamic networks for these two labels have similar distributions across the clusters. Figure A.3 shows the Jaccard distances for all pairs of labels when $K=3$. We also consider integrating the Jaccard distances over a range of $K$. We show the Jaccard distances for all pairs of labels when we tune $K$ from $1$ to $K_{\max}=10$ in Panel (a) of Figure A.4. We then obtain their AUCs and also normalize it by dividing the AUC by $(K_{\max}-1)$ in Panel (b) of Figure A.4. A pair of labels $(i,j)$ with a small normalized AUC of the Jaccard distance suggests that this pair of labels always has similar distributions across the clusters, further implying that the dynamic networks corresponding to these two edges have similar distributions. From Panel (b) of Figure A.4, we observe that edges $1$ and $7$, as well as edges $5$ and $9$, exhibit the most similar functions, whereas edge $6$ is distinct from all other edges.
@@
Figure A.3: Jaccard distances between the distributions across the clusters for all pairs of labels for $K=3$.
(a) Jaccard distances
(b) Normalized AUCs
@@
Figure A.4: Panel (a) shows the Jaccard distances for all pairs of labels when we tune $K$ from $1$ to $K_{\max}=10$. The dashed line represents the maximum value, 11, that the Jaccard distance can achieve. We have total $\binom{13}{2}=78$ lines in Panel (a) to represent all pairs of these $13$ labels in different colors. Panel (b) shows the normalized AUCs for all pairs of labels.
A.2 Additional results for trade dynamic networks
To complement the hierarchical clustering result in Figure 7 of the main text, we also apply CMDS to visualize the distance relationships between products based on $\mathbf{D}^{\star}$, as shown in Figure A.5.
Figure A.5: 22-dimensional embeddings based on $\mathbf{D}^{\star}$ for product trading patterns.
Appendix B Proof of Main Results
B.1 Proof of Theorem 1
@@ B.3 Proof of Theorem 3
$\mathbf{W}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\mathbf{X}_{t_{1}}\mathbf{O}-\mathbf{X}_{t_{2}}\|_{F},\quad\hat{\mathbf{W}}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}_{t_{1}}\mathbf{O}-\hat{\mathbf{X}}_{t_{2}}\|_{F},$
$\mathbf{W}^{(t_{1})}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}_{t_{1}}\mathbf{O}-\mathbf{X}_{t_{1}}\|_{F},\quad\mathbf{W}^{(t_{2})}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}_{t_{2}}\mathbf{O}-\mathbf{X}_{t_{2}}\|_{F}.$
@@
$n^{1/2}|\hat{\mathbf{D}}_{t_{1},t_{2}}-\mathbf{D}_{t_{1},t_{2}}|$ $=\big|\|\hat{\mathbf{X}}_{t_{1}}\hat{\mathbf{W}}-\hat{\mathbf{X}}_{t_{2}}\|_{F}-\|\mathbf{X}_{t_{1}}\mathbf{W}-\mathbf{X}_{t_{2}}\|_{F}\big|$ (B.1)
@@
For $\hat{\mathbf{X}}_{t_{1}}\mathbf{W}^{(t_{1})}-\mathbf{X}_{t_{1}}$ and $\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{2}}$, according to Theorem 2.1 in Tang and Priebe (2018) under the assumption $n\rho_{n}=\Omega(\log n)$ we have
$\|\hat{\mathbf{X}}_{t_{1}}\mathbf{W}^{(t_{1})}-\mathbf{X}_{t_{1}}\|_{F}\lesssim 1\text{ and }\|\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{2}}\|_{F}\lesssim 1$ (B.2)
@@
$\|\mathbf{X}_{t_{1}}\|\asymp(n\rho_{n})^{1/2}\text{ and }\|\mathbf{X}_{t_{2}}\|\asymp(n\rho_{n})^{1/2}.$ (B.3)
@@
$\|\mathbf{W}-\mathbf{W}^{(t_{1})\top}\hat{\mathbf{W}}\mathbf{W}^{(t_{2})}\|\leq\frac{2\|\mathbf{W}^{(t_{1})\top}\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}\|}{\sigma_{\min}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})}.$ (B.4)
Indeed, $\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}$ is invertible. Now suppose $\|\mathbf{W}^{(t_{1})\top}\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}\|<\sigma_{\min}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})$. Then $\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}$ is also invertible and hence, by Theorem 1 in Li (1995) we have
@@
Otherwise if $\|\mathbf{W}^{(t_{1})\top}\hat{\mathbf{X}}_{t_{1}}^{\top}\hat{\mathbf{X}}_{t_{2}}\mathbf{W}^{(t_{2})}-\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}}\|\geq\sigma_{\min}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})$ then, as $\|\mathbf{W}-\mathbf{W}^{(t_{1})\top}\hat{\mathbf{W}}\mathbf{W}^{(t_{2})}\|\leq 2$, Eq. (B.4) holds trivially. By Eq. (B.2) and Eq. (B.3) we have
@@
with high probability. Notice we have $\lambda_{k}(\mathbf{P}_{t_{1}})\asymp n\rho_{n}$ and $\lambda_{k}(\mathbf{P}_{t_{2}})\asymp n\rho_{n}$ for any $k\in[d]$, it follows that $\sigma_{k}(\mathbf{X}_{t_{1}}^{\top}\mathbf{X}_{t_{2}})\asymp n\rho_{n}$ for any $k\in[d]$. Then by Eq. (B.4) we have
$\|\mathbf{W}-\mathbf{W}^{(t_{1})\top}\hat{\mathbf{W}}\mathbf{W}^{(t_{2})}\|\lesssim\frac{(n\rho_{n})^{1/2}}{n\rho_{n}}\lesssim(n\rho_{n})^{-1/2}$ (B.5)
with high probability. Combining Eq. (B.1), Eq. (B.2), Eq. (B.3) and Eq. (B.5) we have